The length of the spring becomes 29cm when a load of 6kg is hung on it. If the elastic potential energy stored in the spring is 2.7J, calculate the original length of the spring?


1
Expert's answer
2021-08-11T09:53:56-0400

Express the spring constant from energy:


"E=\frac12 k\Delta x^2,\\\space\\
E=\frac12 k \Delta x^2,\\\space\\
k=\frac{2E}{\Delta x^2}."

By Hooke's law:


"mg=k\Delta x,\\\space\\
k=\frac{mg}{\Delta x}."

Substitute:


"\frac{mg}{\Delta x}=\frac{2E}{\Delta x^2},\\\space\\
\Delta x=\frac{2E}{mg}=9.2\text{ cm}."

What is "\Delta x"?


"\Delta x=x_f-x_i,\\\space\\
x_i=x_f-\Delta x=29-9.2=19.8\text{ cm}."